OpenStudy (anonymous):

What is the value of a? -2(a-8) + 6a = 36

OpenStudy (dayakar):

remove the parenthesis means bracket first

OpenStudy (dayakar):

can u do that

OpenStudy (anonymous):

I can its just confusing for me.

OpenStudy (dayakar):

multiply with -2 all the terms within the bracket

OpenStudy (anonymous):

16?

OpenStudy (dayakar):

-2a+16+6a=36

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Simplifying -2(a + -8) + 6a = 36 Reorder the terms: -2(-8 + a) + 6a = 36 (-8 * -2 + a * -2) + 6a = 36 (16 + -2a) + 6a = 36 Combine like terms: -2a + 6a = 4a 16 + 4a = 36 Solving 16 + 4a = 36 Solving for variable 'a'. Move all terms containing a to the left, all other terms to the right. Add '-16' to each side of the equation. 16 + -16 + 4a = 36 + -16 Combine like terms: 16 + -16 = 0 0 + 4a = 36 + -16 4a = 36 + -16 Combine like terms: 36 + -16 = 20 4a = 20 Divide each side by '4'. a = 5 Simplifying a = 5

OpenStudy (anonymous):

So a = 5

OpenStudy (dayakar):

above answer is right ,but try to do with ur own

OpenStudy (anonymous):

yes A=5